A poster is to have an area of 630 cm2 with 2.5 cm margins at the bottom and sides and a 5 cm margin at the top. Find the exact dimensions (in cm) that will give the largest printed area.



width : cm



height: cm

1
Expert's answer
2022-05-24T09:54:11-0400

Let "x=" width of the poster, let "y=" height of the poster.

Then


"xy=630"

The largest printed area will be


"A=(x-2.5-2.5)(y-2.5-5)"

Substitute

"A=A(x)=(x-5)(\dfrac{630}{x}-7.5)"

"=630-7.5x-\dfrac{3150}{x}+37.5"

"=667.5-7.5x-\dfrac{3150}{x}, 5\le x\le 84"

Differentiate wih respect to "x"


"A'=-7.5+\dfrac{3150}{x^2}"

Find the critical number(s)


"A'=0=>-7.5+\dfrac{3150}{x^2}=0"

"x=\pm\sqrt{420}"

Since "5\le x\le 84," we take "x=\sqrt{420}"

If "5\le x<\sqrt{420}, A'>0, A" increases.

If "\sqrt{420}<x\le\sqrt{420}, A'<0, A" decreases.

The function "A" has the absolute maximum for "5\le x\le 84" at "x=\sqrt{420}."


"y=\dfrac{630}{\sqrt{420}}=\sqrt{945}"

width : "\sqrt{420}" cm


height: "\sqrt{945}" cm